OpenStudy (anonymous):

The temperature in International Falls, Minnesota, is 9 times colder than the temperature in Cleveland. If the temperature in Cleveland is –2°F, what is the temperature in International Falls? A. –11°F B. –18°F C. –27°F D. 18°F

OpenStudy (ranga):

9 times colder than Cleveland means multiply Cleveland's temperature by 9.

OpenStudy (anonymous):

ok

OpenStudy (ranga):

What is the answer?

OpenStudy (anonymous):

-18

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

i have another

OpenStudy (anonymous):

14.2 • (–6) ÷ 15 • (–4) ÷ 0.4 A. 56.8 B. –3.55 C. –56.8 D. 3.55

OpenStudy (ranga):

\[ 14.2 \times \frac{(-6)}{15} \times \frac{(-4)}{0.4} = ? \]